OpenStudy (anonymous):

Solve for x. -8(-x - 1)+7=2

OpenStudy (luigi0210):

Hello @alpizar_alpizar Welcome to openstudy :) The first thing you wanna do is distribute

OpenStudy (anonymous):

I already did, but its way off :/

OpenStudy (anonymous):

What did you get when you distributed?

OpenStudy (anonymous):

8x+8+7=2

OpenStudy (anonymous):

THat's good. What was you next step?

OpenStudy (anonymous):

I subtracted 7 on both sides. 2-7= -5 and then I brought the 8x-8 down 8x-8=-5 then I added 8 on both sides -5+8=3 Then i divided 8x and 3 by 8